a block of mass m rest on the plank b of mass 3m which is free to slide on a frictionless horizontal surface the coefficient of friction between the block and plank is 0.2.If a horizontal force of magnitude 2 mg is applied to the plank b the accelaration of a relative to the ground respectively ,are

Grade:11

1 Answers

Arun
5 years ago
The weight of a is pressing down on the plank. Let this be W = mg 
Force accelerating a = 0.2 ͯ W = 0.2 ͯ m ͯ g so acceleration of a = 0.2 ͯ g 
Force accelerating plank = (2-0.2) ͯ m ͯ g = 1.8 ͯ m ͯ g 
So acceleration of plank = 1.8 ͯ g/3 = 0.6 ͯ g 
Acceleration of a relative to plank = (0.2 – 0.6)g = -0.4g